Determine the vertex of the graph of the quadratic function f(x)=x2 +4
Answer by harpazo(655)

f(x)= x^2 + 4
x = -b/2a
Let b = 4 and a = 1 from the given function.

x = -4/2(2)
x = -4/4
x = -1
Plug x = -1 into the given function.
f(-1) = (-1)^2 + 4
f(-1) = 1 + 4
f(-1) = 5
Vertex is (-1, 5).